# Robustness of topological phases on aperiodic lattices

### v1

@@ 4.1. Position spectral triples detect strong topological phases
The results of Theorem 3.9 and Theorem 3.32 allows us to compare the groupoid model and the coarse-geometric model on the level of both C*-algebras and K-theory (index pairing). The main theorem is the following:
@@
[…]
@@
  • $\pi_{\omega}^{N}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ is the entrywise extension of the localised regular representation at $\omega\in\Omega_{0}$ as in (3.7);
@@
  • all arrows are given by taking (unbounded) Kasparov products.
@@
We claim that $\pi^{N}_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ maps into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ defined by the standard ample $\omega$-module $\ell^{2}(\omega,\mathcal{K})$. Whenever this holds, then it follows that the K-homology class represented by the spectral triple (3.10)
$\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
equals the pullback along $\pi^{N}_{\omega}$ of the K-homology class represented by (3.25):
$\xi^{\mathrm{Roe}}_{\omega}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\omega)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right).$
We have shown in Theorem 3.9 that $\xi^{\mathrm{Gpd}}_{\omega,N}$ is a locally compact position spectral triple. This implies that $\pi_{\omega}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R}))\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ by Theorem 3.23. Since $\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$ is dense in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$, we conclude that the image of $\pi_{\omega}^{N}$ is contained in $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$. Then it follows from the functoriality of the Kasparov product in Lemma 2.13 that the bottom left triangle commutes.
The commutativity of the top right triangle follows from the construction of $\xi^{\mathrm{Gpd}}_{\omega}$ in Theorem 3.9. Therefore, we conclude that the entire diagram commutes. ∎
@@
Consider the periodic square lattice $\Lambda=\mathbb{Z}^{d}$. Then $\Omega_{\Lambda}\simeq\mathbb{T}^{d}$ and $\Omega_{0}=\{\omega\}$ is a singleton, in which $\omega$ can be chosen to be any point in $\Omega_{\Lambda}$ (cf. [Bourne-Mesland:Aperiodic]*Example 2.7). There is a unique evaluation point $\operatorname{ev}_{\omega}\colon\{\omega\}\to\mathbb{C}$, which gives an isomorphism $\operatorname{K}_{0}(\mathbb{C})\xrightarrow{\sim}\mathbb{Z}$. The spectral triple $\xi^{\mathrm{Gpd}}_{\omega,N}$ now becomes a spectral triple over $\mathrm{C}^{*}(\{\omega\}\rtimes\mathbb{Z}^{d})\otimes\mathbb{M}_{N}(\mathbb{R})\simeq\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d}))$, given by the exterior product of (3.29) with the imprimivitity bimodule $(\mathbb{M}_{N}(\mathbb{R}),\mathcal{K},0)$ induced by any rank-$N$ corner embedding $\mathbb{R}^{N}\hookrightarrow\mathcal{K}$. Thus we have recovered the commutative diagram (3.28) in [Ewert-Meyer:Coarse_geometry]*Theorem 7.
Example 4.4 (Kubota’s construction).
If $\Lambda$ is aperiodic, i.e. there is no non-zero $x\in\mathbb{R}^{d}$ such that $\Lambda=\Lambda-x$, then Kubota has constructed in [Kubota:Controlled_topological_phases]*Lemma 2.24, a surjective groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\Lambda}\to\mathcal{G}_{\Lambda}$, from the coarse groupoid $\mathcal{G}^{\mathrm{coa}}_{\Lambda}$ (cf. [Skandalis-Tu-Yu:Coarse_groupoids]) to the groupoid of Delone sets, which induces an inclusion $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$ into the uniform Roe C*-algebra of $\Lambda$, and hence into $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ because $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$. We claim that his construction indeed works for all $\omega\in\Omega_{0}$, and coincides with the representation $\pi_{\omega}$ as in (3.6).
Let $\Lambda$ be an aperiodic Delone set. Then there exists an open, dense embedding $\varphi_{\Lambda}\colon\Lambda\hookrightarrow\Omega_{0}$ sending $x$ to $\Lambda-x$. Therefore, $\Omega_{0}$ is a compactification of $\Lambda$. It follows from the universal property of the Stone–Čech compactification that there exists a unique surjection $\overline{\varphi}_{\Lambda}\colon\beta\Lambda\to\Omega_{0}$ extending $\varphi_{\Lambda}$.
The coarse groupoid $\mathcal{G}^{\mathrm{coa}}_{\Lambda}$ is the union of all $\overline{\mathcal{E}}$, where $\mathcal{E}$ is a controlled subset of $\Lambda\times\Lambda$ (cf. Definition 3.13), and the closure is taken inside the Stone–Čech compactification of $\Lambda\times\Lambda$. Let $\mathcal{E}\subseteq\Lambda\times\Lambda$ be controlled, then the function
$\mathcal{E}\to\mathbb{R}^{d},\quad(x,y)\mapsto y-x$
is bounded, hence extends to $\beta\mathcal{E}$. Therefore, there is a surjective groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\Lambda}\to\mathcal{G}_{\Lambda}$, which uniquely extends the map
$\Lambda\times\Lambda\to\mathcal{G}_{\Lambda},\quad(x,y)\mapsto(\varphi_{\Lambda}(x),y-x).$
This induces an injective *-homomorphism
$\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda).$
Kubota’s construction holds for all $\omega\in\Omega_{0}$, disregarding whether or not it is aperiodic. (Even if $\Lambda$ is aperiodic, an element $\omega\in\Omega_{0}$ can still be periodic for some $x\in\mathbb{R}^{d}$.) To see this, note that
$\varphi_{\omega}\colon\omega\to\Omega_{0},\qquad x\mapsto\omega-x$
is a continuous map into a compact Hausdorff space. It fails to be injective if $\Lambda$ is not aperiodic, but still extends uniquely to a continuous map $\overline{\varphi}_{\omega}\colon\beta\omega\to\Omega_{0}$. Thus Kubota’s construction gives a (not necessarily surjective) groupoid homomorphism $\mathcal{G}_{\omega}^{\mathrm{coa}}\to\mathcal{G}_{\Lambda}$, and hence a *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}^{\mathrm{coa}}_{\omega})\simeq\mathrm{C}^{*}_{\mathrm{u,Roe}}(\omega)$. Composing with a rank-one corner embedding yields a map into $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\omega)\otimes\mathbb{K}\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$.
On the other hand, the *-homomorphism $\pi^{1}_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ in (3.6) maps $f\in\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ to the operator $(\pi_{\Lambda}(f)_{x,y})_{x,y\in\Lambda}$, whose matrix elements are given by
$\pi_{\omega}(f)_{x,y}=\left\langle x\,\middle|\pi_{\omega}(f)\middle|\,\mathopen{}y\right\rangle=f(\omega-x,y-x)=f(\varphi_{\omega}(x),y-x).$
Therefore, it is induced by Kubota’s groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\omega}\to\mathcal{G}_{\Lambda}$.
@@ 4.2. Stacked topological phases are weak
We shall explain in this section how this observation can be generalised to aperiodic lattices and higher-dimensional cases. That is, we consider Delone sets of the form $\Lambda\times L$, where $L$ is another Delone set. If $L$ has dimension $1$, then we may think of $\Lambda\times L$ as “stacking” $\Lambda$ along the direction of $L$. This also gives a *-homomorphism groupoid *-homomorphism $\varphi^{\mathrm{Gpd}}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$, which induces maps
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$
for all $k$. Such maps can be interpreted as “stacking” topological phases living on $\Lambda$ along the direction $L$. We will show that such topological phases are always weak, in the sense that they vanish in the K-theory of Roe C*-algebras.
We fix some notations. Let $\Lambda\subseteq\mathbb{R}^{m}$ be an $(r_{\Lambda},R_{\Lambda})$-Delone set and $L\subseteq\mathbb{R}^{n}$ be an $(r_{L},R_{L})$-Delone set. We write $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}$ and $\Omega_{L}$ for the closure of the orbit of $\Lambda\times L$, $\Lambda$ and $L$. To distinguish, we write
[…]
@@
$\Lambda\times L\mathrel{\vcentcolon=}\left\{(x,a)\;\middle|\;x\in\Lambda,a\in L\right\}\subseteq\mathbb{R}^{m+n}$
@@
#$\#(\operatorname{B}((x,a),r)\cap(\Lambda\times L))$$\leq\#((\operatorname{B}(x,r_{\Lambda})\times\operatorname{B}(a,r_{L}))\cap(\Lambda\times L))$
=#$=\#((\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,r_{L})\cap L))$
=#$=\#(\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,r_{L})\cap L)$
@@
#$\#(\operatorname{B}((x,a),R)\cap(\Lambda\times L))$$\geq\#((\operatorname{B}(x,R_{\Lambda})\times\operatorname{B}(a,R_{L}))\cap(\Lambda\times L))$
=#$=\#((\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,R_{L})\cap L))$
=#$=\#(\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,R_{L})\cap L)$
@@
A convenient way to define the “stacking” map is by describing the C*-algebra of the product Delone set as a minimal tensor product. We note the following well-known lemma:
@@
$\Phi\colon\mathrm{C}^{*}(\mathcal{G}_{1})\otimes\mathrm{C}^{*}(\mathcal{G}_{2})\xrightarrow{\sim}\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2}),\quad\Phi(f_{1}\otimes f_{2})(\gamma_{1},\gamma_{2})\mathrel{\vcentcolon=}f_{1}(\gamma_{1})f_{2}(\gamma_{2}).$
where $\otimes$ refers to the minimal tensor product.
@@
By Lemma 4.6, it suffices to prove that the étale groupoids $\mathcal{G}_{\Lambda\times L}$ and $\mathcal{G}_{\Lambda}\times\mathcal{G}_{L}$ are isomorphic.
@@
The translation action of $\mathbb{R}^{m+n}=\mathbb{R}^{m}\times\mathbb{R}^{n}$ preserves $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$. By Proposition 3.1, both $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$ are closed, hence their product is a closed subset of $\mathcal{M}(\mathbb{R}^{m+n})$. Since the product of the orbits of $\Lambda$ and $L$ is contained in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$, we conclude that $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}\times\Omega_{L}$ are contained in it. Hence, $\Omega_{\Lambda\times L}\subseteq\Omega_{\Lambda}\times\Omega_{L}$ as it is the smallest closed subset in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$, which contains the product of the joint orbits of $\Lambda$ and $L$.
We claim that $\Omega_{\Lambda\times L}\supseteq\Omega_{\Lambda}\times\Omega_{L}$ as well. Let $(\omega,\ell)\in\Omega_{\Lambda}\times\Omega_{L}$, then there exists nets $(x_{\alpha})_{\alpha\in A}\subseteq\mathbb{R}^{m}$ and $(a_{\beta})_{\beta\in B}\subseteq\mathbb{R}^{n}$ such that
[…]
@@
[…]
Then we conclude that $\Omega_{\Lambda\times L}\simeq\Omega_{\Lambda}\times\Omega_{L}$ as $\mathbb{R}^{m+n}$-spaces.
As a consequence, the action groupoids $\Omega_{\Lambda\times L}\rtimes\mathbb{R}^{m+n}$ and $(\Omega_{\Lambda}\times\Omega_{L})\rtimes\mathbb{R}^{m+n}$ are isomorphic topological groupoids, and have homeomorphic abstract transversals $\Omega_{0}^{\Lambda\times L}$ and $\Omega_{0}^{\Lambda}\times\Omega_{0}^{L}$. This implies an isomorphism of étale groupoids

### v2

@@ 4.1. Position spectral triples detect strong topological phases
The results of Theorem˜3.12 and Theorem˜3.33 allows us to compare the groupoid model and the coarse-geometric model on the level of both C*-algebras and K-theory (index pairing). The main theorem is the following:
@@
(4.2) $\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))$$\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathrm{C}(\Omega_{0}))$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))$$\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R}).$$\scriptstyle{\prescript{}{d}{\lambda}_{\Omega_{0}}}$$\scriptstyle{\pi^{N}_{\omega}}$$\scriptstyle{\xi^{\mathrm{Gpd}}_{\omega,N}}$$\scriptstyle{\operatorname{ev}_{\omega}}$$\scriptstyle{\xi^{\mathrm{Roe}}_{\omega}}\scriptstyle{\sim}$
@@
  • $\pi_{\omega}^{N}\mathrel{\vcentcolon=}\pi_{\omega}\otimes e_{N}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega)\otimes\mathcal{K})$ is the entrywise extension of the localised regular representation at $\omega\in\Omega_{0}$ as in (3.10);
@@
  • all arrows are given by taking Kasparov products.
@@
We claim that $\pi^{N}_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega)\otimes\mathcal{K})$ maps into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ defined by the standard ample $\omega$-module $\ell^{2}(\omega,\mathcal{K})$. Whenever this holds, then it follows that the KK-class represented by the spectral triple (3.13)
$\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)$
equals the pullback along $\pi^{N}_{\omega}$ of the KK-class represented by (3.28):
$\xi^{\mathrm{Roe}}_{\omega}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\omega)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right).$
We have shown in Theorem˜3.12 that $\xi^{\mathrm{Gpd}}_{\omega,N}$ is a position spectral triple. This implies that $\pi_{\omega}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R}))\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ by Theorem˜3.26. Since $\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$ is dense in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$, the image of $\pi_{\omega}^{N}$ is contained in $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$. Then it follows from the functoriality of the Kasparov product in Lemma˜2.10 that the bottom left triangle commutes.
The commutativity of the top right triangle follows from the construction of $\xi^{\mathrm{Gpd}}_{\omega}$ in Theorem˜3.12. Therefore, we conclude that the entire diagram commutes. ∎
@@
Consider the periodic square lattice $\Lambda=\mathbb{Z}^{d}$. Then $\Omega_{\Lambda}\simeq\mathbb{T}^{d}$ and $\Omega_{0}=\{\omega\}$ is a singleton, in which $\omega$ can be chosen to be any point in $\Omega_{\Lambda}$ (cf. [Bourne-Mesland:Topological_phases]*Example 2.7). There is a unique evaluation point $\operatorname{ev}_{\omega}\colon\{\omega\}\to\mathbb{C}$, which gives an isomorphism $\operatorname{KO}^{0}(\operatorname{pt})=\operatorname{KO}_{0}(\mathbb{R})\xrightarrow{\sim}\mathbb{Z}$. The spectral triple $\xi^{\mathrm{Gpd}}_{\omega,N}$ is nothing but the position spectral triple $\xi_{\mathbb{Z}^{d},N}$ over the real group C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$ in (2.14). And the diagram (4.2) recovers the commutative diagram in [Ewert-Meyer:Coarse_geometry]*Theorem 7.
@@ 4.2. Stacked topological phases are weak
We shall explain in this section how this observation can be generalised to aperiodic lattices and higher-dimensional cases. That is, we consider Delone sets of the form $\Lambda\times L$, where $L$ is another Delone set. If $L$ has dimension $1$, then we may think of $\Lambda\times L$ as “stacking” $\Lambda$ along the direction of $L$ (cf. Figure˜4.1). This also gives a *-homomorphism groupoid *-homomorphism $\varphi^{\mathrm{Gpd}}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$, which induces maps
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$
for all $n$. Such maps can be interpreted as “stacking” topological phases living on $\Lambda$ along the direction $L$. We will show that such topological phases are always weak, in the sense that they vanish in the K-theory of Roe C*-algebras.
xxyyzz
Figure 4.1. “Stacked” topological phases along the $z$-direction
We fix the following notation. Let $\Lambda\subseteq\mathbb{R}^{p}$ be an $(r_{\Lambda},R_{\Lambda})$-Delone set and $L\subseteq\mathbb{R}^{q}$ be an $(r_{L},R_{L})$-Delone set. We write $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}$ and $\Omega_{L}$ for the closure of the orbit of $\Lambda\times L$, $\Lambda$ and $L$. To distinguish, we write
$\Omega^{\Lambda\times L}_{0}\mathrel{\vcentcolon=}\{\mu\in\Omega_{\Lambda\times L}\mid 0\in\mu\},\quad\Omega^{\Lambda}_{0}\mathrel{\vcentcolon=}\{\omega\in\Omega_{\Lambda}\mid 0\in\omega\},\quad\Omega^{L}_{0}\mathrel{\vcentcolon=}\left\{\ell\in\Omega_{L}\;\middle|\;0\in\ell\right\}$
@@
$\Lambda\times L\mathrel{\vcentcolon=}\left\{(x,a)\;\middle|\;x\in\Lambda,a\in L\right\}\subseteq\mathbb{R}^{p+q}$
@@
$\#(\operatorname{B}((x,a),r)\cap(\Lambda\times L))$ $\leq\#((\operatorname{B}(x,r_{\Lambda})\times\operatorname{B}(a,r_{L}))\cap(\Lambda\times L))$
$=\#((\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,r_{L})\cap L))$
$=\#(\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,r_{L})\cap L)$
@@
$\#(\operatorname{B}((x,a),R)\cap(\Lambda\times L))$ $\geq\#((\operatorname{B}(x,R_{\Lambda})\times\operatorname{B}(a,R_{L}))\cap(\Lambda\times L))$
$=\#((\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,R_{L})\cap L))$
$=\#(\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,R_{L})\cap L)$
@@
A convenient way to define the “stacking” map is by describing the C*-algebra of the product Delone set as a spatial tensor product. We note the following well-known lemma:
@@
$\Phi\colon\mathrm{C}^{*}(\mathcal{G}_{1})\otimes\mathrm{C}^{*}(\mathcal{G}_{2})\xrightarrow{\sim}\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2}),\quad\Phi(f_{1}\otimes f_{2})(\gamma_{1},\gamma_{2})\mathrel{\vcentcolon=}f_{1}(\gamma_{1})f_{2}(\gamma_{2}),$
where $\otimes$ refers to the spatial tensor product.
Proof.
This lemma is well-known, and we sketch a proof here.
Write $L^{2}(\mathcal{G}_{1})$, $L^{2}(\mathcal{G}_{2})$ and $L^{2}(\mathcal{G}_{1}\times\mathcal{G}_{2})$ for the Hilbert C*-modules over $\mathrm{C}_{0}(\Omega_{1})$, $\mathrm{C}_{0}(\Omega_{2})$ and $\mathrm{C}_{0}(\Omega_{1}\times\Omega_{2})$ as in Section˜3.1.1. We claim that $L^{2}(\mathcal{G}_{1}\times\mathcal{G}_{2})$ is canonically isomorphic to the external tensor product Hilbert C*-module $L^{2}(\mathcal{G}_{1})\otimes L^{2}(\mathcal{G}_{2})$ under the identification
$\Psi\colon\mathrm{C}_{0}(\Omega_{1})\otimes\mathrm{C}_{0}(\Omega_{2})\simeq\mathrm{C}_{0}(\Omega_{1}\times\Omega_{2}),\quad\Psi(f\otimes g)(x,y)\mathrel{\vcentcolon=}f(x)\cdot g(y).$
To see this, let $f_{1},f_{2}\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{1})$ and $g_{1},g_{2}\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{2})$, Then their tensor products $f_{1}\otimes g_{1},f_{2}\otimes g_{2}$ belong to $\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{1}\times\mathcal{G}_{2})$. Given $x\in\Omega_{1}$ and $y\in\Omega_{2}$, we have
$\left\langle f_{1},f_{2}\right\rangle(x)=\sum_{\gamma\in r^{-1}(x)}f_{1}(\gamma)f_{2}(\gamma),\quad\left\langle g_{1},g_{2}\right\rangle(y)=\sum_{\eta\in r^{-1}(y)}g_{1}(\eta)g_{2}(\eta),$
and
$\left\langle f_{1}\otimes g_{1},f_{2}\otimes g_{2}\right\rangle(x,y)$ $=\sum_{(\gamma,\eta)\in r^{-1}(x,y)}(f_{1}\otimes g_{1})(\gamma,\eta)(f_{2}\otimes g_{2})(\gamma,\eta)$
$=\sum_{\gamma\in r^{-1}(x)}\sum_{\eta\in r^{-1}(y)}f_{1}(\gamma)g_{2}(\gamma)g_{1}(\eta)g_{2}(\eta)$
$=\langle f_{1},f_{2}\rangle(x)\cdot\langle g_{1},g_{2}\rangle(y).$
[…]
$\Phi\colon\mathbb{B}_{\mathrm{C}_{0}(\Omega_{1})}(L^{2}(\mathcal{G}_{1}))\otimes\mathbb{B}_{\mathrm{C}_{0}(\Omega_{2})}(L^{2}(\mathcal{G}_{2}))\to\mathbb{B}_{\mathrm{C}_{0}(\Omega_{1}\times\Omega_{2})}(L^{2}(\mathcal{G}_{1}\times\mathcal{G}_{2})).$
In particular, for $f_{1}\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{1})$ and $f_{2}\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{2})$, viewed as convolution operators on $L^{2}(\mathcal{G}_{1})$ and $L^{2}(\mathcal{G}_{2})$ via their regular representations $\pi_{i}\colon\mathrm{C}^{*}(\mathcal{G}_{i})\to\mathbb{B}_{\mathrm{C}_{0}(\Omega_{i})}(\mathcal{G}_{i})$ for $i=1,2$, then $\Phi$ sends $f_{1}\otimes f_{2}$ to the convolution operator $f_{1}\otimes f_{2}\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{1}\times\mathcal{G}_{2})$ on $L^{2}(\mathcal{G}_{1}\times\mathcal{G}_{2})$. Passing to the norm closure, we find that $\Phi$ maps $\mathrm{C}^{*}(\mathcal{G}_{1})\otimes\mathrm{C}^{*}(\mathcal{G}_{2})$ into $\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2})$, and has dense range as $\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{1}\times\mathcal{G}_{2})$ is dense in $\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2})$. Thus $\Phi$ is both injective and has dense range, hence a *-isomorphism
$\mathrm{C}^{*}(\mathcal{G}_{1})\otimes\mathrm{C}^{*}(\mathcal{G}_{2})\xrightarrow{\sim}\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2}).\qed$
@@
By Lemma˜4.5, it suffices to prove that the étale groupoids $\mathcal{G}_{\Lambda\times L}$ and $\mathcal{G}_{\Lambda}\times\mathcal{G}_{L}$ are isomorphic.
@@
The translation action of $\mathbb{R}^{p+q}=\mathbb{R}^{p}\times\mathbb{R}^{q}$ preserves $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{p})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{q})$. By Proposition˜3.4, both $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$ are closed, hence their product is a closed subset of $\mathcal{M}(\mathbb{R}^{p+q})$. Since the product of the orbits of $\Lambda$ and $L$ is contained in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{p})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{q})$, we conclude that $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}\times\Omega_{L}$ are contained in it. Hence, $\Omega_{\Lambda\times L}\subseteq\Omega_{\Lambda}\times\Omega_{L}$ as it is the smallest closed subset in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{p})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{q})$, which contains the product of the joint orbits of $\Lambda$ and $L$.
We claim that $\Omega_{\Lambda\times L}\supseteq\Omega_{\Lambda}\times\Omega_{L}$ as well. Let $(\omega,\ell)\in\Omega_{\Lambda}\times\Omega_{L}$, then there exists nets $(x_{\alpha})_{\alpha\in A}\subseteq\mathbb{R}^{p}$ and $(a_{\beta})_{\beta\in B}\subseteq\mathbb{R}^{q}$ such that
$\Lambda+x_{\alpha}\to\omega\quad\text{and}\quad L+a_{\beta}\to\ell\quad\text{in the weak{}{*}-topology.}$
@@
$(\Lambda+x_{\alpha},L+a_{\beta})_{(\alpha,\beta)\in A\times B}\to(\omega,\ell)\quad\text{in the weak{}{*}-topology}.$
Then we conclude that $\Omega_{\Lambda\times L}\simeq\Omega_{\Lambda}\times\Omega_{L}$ as $\mathbb{R}^{p+q}$-spaces.
As a consequence, the action groupoids $\Omega_{\Lambda\times L}\rtimes\mathbb{R}^{p+q}$ and $(\Omega_{\Lambda}\times\Omega_{L})\rtimes\mathbb{R}^{p+q}$ are isomorphic topological groupoids, and have homeomorphic abstract transversals $\Omega_{0}^{\Lambda\times L}$ and $\Omega_{0}^{\Lambda}\times\Omega_{0}^{L}$. This implies an isomorphism of étale groupoids